OpenStudy (anonymous):

how do i find the slope between the pairs of points (-6, 4) and (0,0)

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = (0-4)/(0-(-6)) m = (0-4)/(0+6) I'll let you finish up. Don't forget to fully simplify and reduce.